st+5b-bs-5t how do you factor out this problem?
Found 2 solutions by drj, RAY100:
Answer by drj(1380)

st+5b-bs-5t how do you factor out this problem?

Step 1. Rearrange the expression

st+5b-bs-5t=st-bs-5t+5b

Step 2. Factor out s in the first two terms and factor our -5 in the last two terms.

st-bs-5t+5b=s(t-b)-5(t-b)

Step 3. Note that (t-b) is common in the parenthesis...so factor it out

s(t-b)-5(t-b)=(t-b)(s-5)

Step 4. ANSWER: st+5b-bs-5t=(t-b)(s-5)

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.


Answer by RAY100(1637)

Usually when we see 4 different terms we try " grouping"
.
st +5b -bs -5t,,,rearrange terms
.
st - bs +5b-5t
.
s(t-b) +5(b-t),,,,change signs 2nd term
.
s(t-b) -5(t-b),,,,,,,,,,,,,,now comes hard part,,,,,treat this as s(x) -5(x),,where
.
x = (t-b),,,,then x(s-5) is factored form,,,subst back in for x
.
,,,,,,(t-b)(s-5)
.
(t-b)(s-5),,,,,answer
.